OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x))

jimthompson5910 (jim_thompson5910):

What you do is plug in g(x) into the f(x) function, then simplify f(x) = -4x + 7 f(g(x)) = -4( g(x) ) + 7 ... replace each x with g(x) f(g(x)) = -4( 10x-6 ) + 7 ... replace each g(x) on the right side with 10x-6 f(g(x)) = -4*10x-4*( -6 ) + 7 ... distribute f(g(x)) = -40x + 24 + 7 ... multiply f(g(x)) = -40x + 31 ... combine like terms.

OpenStudy (anonymous):

Got it, thankyouu !!

jimthompson5910 (jim_thompson5910):

yw